OpenStudy (anonymous):

what is (1-(4/x))/(1-(2/x)-(8/x^2)) in simplest form? please explain how you get your answer

OpenStudy (anonymous):

also known as:\[\frac{ 1-\frac{ 4 }{ x } }{ 1-\frac{ 2 }{ x }-\frac{ 8 }{ x ^{2} } }\]

OpenStudy (anonymous):

\[\huge\frac{1-\frac{4}{x}}{1-\frac{2}{x}-\frac{8}{x^2}}\]

OpenStudy (anonymous):

same idea as before find a common denominator

OpenStudy (anonymous):

get the answer is one step by multiplying top and bottom by \(x^2\) cancelling merrily as you go

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what satellite73 said works too, and faster

OpenStudy (anonymous):

so...\[\frac{ \frac{ x ^{2} }{ x ^{2} }-\frac{ 4x }{ x ^{2} } }{ \frac{ x ^{2} }{ x ^{2} }-\frac{ 2x }{ x ^{2} }-\frac{ 8 }{ x ^{2} } }\]?

OpenStudy (anonymous):

oh no !!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\huge\frac{(1-\frac{4}{x})}{(1-\frac{2}{x}-\frac{8}{x^2})}\times \frac{x^2}{x^2}\]

OpenStudy (anonymous):

distribute what is \(1\times x^2\) ?

OpenStudy (anonymous):

\[x ^{2}\]

OpenStudy (anonymous):

ok and \(-\frac{4}{x}\times x^2\) ?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

really? don't believe you \[\frac{-4x^2}{x}=?\]?

OpenStudy (anonymous):

im confused im sorry

OpenStudy (anonymous):

cancel an x top and bottom or subtract the exponent either way you get \[-\frac{4}{x}\times x^2=-4x\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so the numerator is \[x^2-4x\] how about the denominator?

OpenStudy (anonymous):

ok \[x ^{2}-2x-8x\]is the denominator?

OpenStudy (anonymous):

close

OpenStudy (anonymous):

\[x^2-2x-8\] the last term is just \(-8\) cause the \(x^2\) cancel

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

so we are at \[\frac{x^2-4x}{x^2-2x-8}\] now i think we can factor top and bottom

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

matter of fact it am sure of it do you see how to factor the top?

OpenStudy (anonymous):

yes it would be x(x-4)

OpenStudy (anonymous):

and the bottom?

OpenStudy (anonymous):

and the bottom would be (x-4)(x+2)

OpenStudy (anonymous):

then cancel and be done!

OpenStudy (anonymous):

then you can cancel out the (x-4)s ok thank you

OpenStudy (anonymous):

yw